Calculate the wavelength of maximum emission for a tungsten lamp operating at 2870 K. By what factor will its output intensity increase if we raise its temperature to 3000 K?
Answer #1

× T = 2.898×10-3 m.K

= 2.898×10-3m.K /2870K = 1.01×10-6 m

Wavelength of maximum emission = 1.01×10-6 m. (answer)

Intensity is directly proportional to 4th power of temperature. (T​​​​​​4 )

I​​​​​​2 = (T2/T1)4 × I​​​​​​1

= (3000/2870)4 × I​​​​​​1

= 1.194 I​​​​​​1 ~ 1.2× I​​​​​​1

The output intensity would increases by a factor of 1.2 times.(answer)